Water is flowing continuously from a tap having an internal diameter 8*10-3m. The water velocity as it leaves the tap is 0.4 ms-1. The diameter of the water stream at a distance 2*10-1m below the tap is close to :
(a) 7.5*10-3m
(b) 9.6*10-3m
(c) 3.6*10-3m
(d) 5.4*10-3m

Grade:11

2 Answers

Kevin Nash
9 years ago

(c). From Bernoulli’s theorem,

P0 + 1/2 pv21pgh = p0 + 1/2 pv22 + 0

V2 = vv21 + 2gh = v0.16 + 2 * 10 *0.2 = 2.03 m/s

From equation of continuity

A2v2 = A1v1

pD22/4 * v2 = pD21/4v1

? D1D2 vv1/v2 = 3.55 * 10-3 m

ankit singh
3 years ago

Given data,

d1=8×103m
v1=0.4ms1
h=0.2m
According to equation of motion,
v2=v12+2gh
=(0.4)2+2×10×0.2
2ms1
According to equation of continuity
 a1v1=a2v2
π×(28×103)2×0.4=π×(2d2)2×2
d2=3.6×103m